Assume the risk-free rate is 6% and the market risk premium is 6%. The stock of Physicians Care Network (PCN) has a beta of 1.5. The last dividend paid by PCN (D0) was $2 per share. What would PCN’s stock value be if the dividend were expected to grow at a constant rate of negative 5%.

Choice: $6.00

Choice: $9.50

Choice: $13.45

Choice: $17.60

Assume the risk-free rate is 6% and the market risk premium is 6%. The stock of Physicians Care Network (PCN) has a beta of 1.5. The last dividend paid by PCN (D0) was $2 per share. What would PCN’s stock value be if the dividend were expected to grow at a constant rate of 0%.

Choice: $0.00

Choice: $5.05

Choice: $13.33

Choice: $20.40

Assume the risk-free rate is 6% and the market risk premium is 6%. The stock of Physicians Care Network (PCN) has a beta of 1.5. The last dividend paid by PCN (D0) was $2 per share. What would PCN’s stock value be if the dividend were expected to grow at a constant rate of 10%.

Choice: $35.00

Choice: $40.00

Choice: $44.00

Choice: $50.00

Assume the risk-free rate is 6% and the market risk premium is 6%. The stock of Physicians Care Network (PCN) has a beta of 1.0. The last dividend paid by PCN (D0) was $2 per share. What would be the stock value if the growth rate were 10%?

Choice: $50.00

Choice: $100.00

Choice: $110.00

Choice: $115.00
Answer #1

First of All here i am explaining the formula as per gordon growth model. i.e

P0= D0(1+g)/(Kc-g)

Where

P0 = stock value

D0= current dividend or divident paid

g = growth rate

Kc = cost of capital

As per CAPM

Kc= Rf+(Rm-Rf)Beta

Where

Rf= risk free rate of return

Rm-Rf = market risk premium

Answer to question no (1)

Ans is $ 9.5

Calculations

Kc= 6+6*1.5 = 15%

P0= 2(1+(-.05))/ (.15-(-.05))

= 2(1-.05)/(.15+.05)

= 2*.95/.20

P0 = 1.9/.20 = $ 9.5

Answer to question no (2)

Ans is $ 13.33

Calculations

Kc= 15% (As per above calculations)

P0 = 2(1+0)/(.15+0)

= 2/.15

= $ 13.33

Answer to question no (3).

Ans is $ 44.00

calculations

Kc= 15%

P0 = 2(1+.10)/(.15-.10)

= 2.2/.05

= $ 44

Answer to question no (4)

Ans is $110

calculations

Kc = 6+6*1.0

= 12%

P0 = 2(1.10)/(.12-.10)

= 2.2/.02

= $ 110

note:- please feel free to ask in comment section if you have any query.

Request , Don't forget to press thumbs up